Dakota Trucking Company (DTC) is evaluating a potential lease for a truck with a 4-year life that costs $40,000 and falls into the MACRS 3-year class. If the firm borrows and buys the truck, the loan rate would be 10%. The truck will be used for 4 years, at the end of which time it will be sold at an estimated residual value of $10,000. If DTC buys the truck, it would purchase a maintenance contract that costs $1,000 per year, payable at the end of each year. The lease terms call for a $10,000 lease payment (4 payments total) at the beginning of each year. DTC's tax rate is 40%. Should the firm lease or buy? (Note: MACRS rates for Years 1 to 4 are 0.33, 0.45, 0.15, and 0.07.) Question 3) What is IPO underpricing?
Answer #1
a) 0 1 2 3 4
LEASING:
After tax lease rent [10000*(1-40%) $         -6,000 $         -6,000 $        -6,000 $        -6,000
After tax cash outflows $         -6,000 $         -6,000 $        -6,000 $        -6,000
b) BUYING:
Loan installment = 40000*0.1*1.1^4/(1.1^4-1) = $        12,619
Loan amortization:
Beginning balance of loan $        40,000 $       31,381 $        21,900 $        11,471
Interest at 10% $           4,000 $          3,138 $          2,190 $          1,147
Total $        44,000 $       34,519 $        24,090 $        12,618
Installment $        12,619 $       12,619 $        12,619 $        12,618
Ending balance of loan $        31,381 $       21,900 $        11,471 $                   0
Depreciation $        13,200 $       18,000 $          6,000 $          2,800
After cash flows of buying:
Principal repayment $         -8,619 $        -9,481 $      -10,429 $      -11,471
After tax interest $         -2,400 $        -1,883 $        -1,314 $            -688
Tax shield on depreciation at 40% $           5,280 $          7,200 $          2,400 $          1,120
After tax residual value = 10000*(1-40%) = $          6,000
After tax maintenance cost (1000*60%) $             -600 $            -600 $            -600 $            -600
After tax cash outflows $         -6,339 $        -4,764 $        -9,943 $         -5,639
c) PVIF at 6% [PVIF = 1/1.06^n] 1 0.94340 0.89000 0.83962 0.79209
PV of leasing $         -6,000 $         -5,660 $        -5,340 $        -5,038
Total PV $      -22,038
PV of buying $         -5,980 $        -4,240 $        -8,348 $         -4,467
Total PV $      -23,035
d) Leasing is preferred as the PV of cash outflows is lower.